Thornton Computer Services, Inc. has been in business for sik months. The following are basic operating data for that period: Month Dec Sept. 260 oct. Service hours 120 136 420 320 330 6,000 6,800 $13,000 $21,000 $16,000 $16, 500 $4,300 $5,300 $ 7,100 $11,200 $ 9,100 s10, 600 Operating costs Required a. What is the averag b. Use the high-low method to estimate the total monthly fixed cost and the variable cost per hour c. Determine the average contribution margin per hour. e service revenue per hour in each month and the overall average for the six-month period? Complete this question by entering your answers in the tabs below Required A Required B Required C What is the average service revenue per hour in each month and the overall average for the six-month period? July Aug Sept. Oct. Nov. Dec. Total Average Required B>
Answer #1

Required A. Calculation of average revenue per hour for Each Month And Overall average for 6 Month :

Average revenue per hour = Revenue/ Service Hours

July August September October November December Total
Revenue 6000 6800 13000 21000 16000 16500 79300
Service Hours 120 136 260 420 320 330 1586
Average service revenue per hour 50 50 50 50 50 50 50

Hence,

For  each month July to December Average service revenue per hour = 50 $ Per Service hour

Overall 6 month Average service revenue per hour = $50 Per Service hour

Required B. Using High Low Method Monthly Fix Cost & Variable Cost calculation :-

Month Operating Cost Service Hours Variable Cost
Oct high 11200 420
July Low 4300 120
Difference 6900 300 (6900/300)        23.00

Fixed Cost = Operating Cost in October - (Variable cost per unit * Service   hours in October)

= 11200 - (420*23.00)

= 11200 -9660 = 1540

Therefore, Monthly Fix Cost = $1540  

Variable Cost Per Hour = $23.00

Required C. Average Contribution margin per hour

Particulars Amount
Average Revenue per hour(Calculated in Part A)    50
(Less)Variable cost per hour 23
Average Contribution margin per hour 27